OpenStudy (anonymous):

Explain why (–4x)0 = 1, but –4x0 = –4.

OpenStudy (agreene):

I assume you're asking why: \[(-4x)^0=1\] whilst \[-4x^0=-4\] and the answer is that anything to the 0 is defined as 1. and the order of operations: (-4x)^0 is saying the variable and coefficient are to the 0 while -4x^0 is only taking the variable to the 0 which means: \[-4x^0\implies-4(1)=-4\]

OpenStudy (anonymous):

what agreen said. for the same reason that \[2\times 3^2=2\times 9=18\] but \[(2\times 3)^2=6^2=36\]